# Statistics-encoded tensor network approach in disordered quantum many-body spin chains

### v1

@@ Abstract
Simulating the dynamics of quantum many-body systems with disorder is a fundamental challenge. In this work, we propose a general approach—the statistics-encoded tensor network (SeTN)—to study such systems. By encoding disorder into an auxiliary layer and averaging separately, SeTN restores translational invariance, enabling a well-defined transfer matrix formulation. We derive a universal criterion, $n\gg\alpha^{2}t^{2}$, linking discretization $n$, disorder strength $\alpha$, and evolution duration $t$. This sets the resolution required for faithful disorder averaging and shows that encoding is most efficient in the weak-disorder, typically chaotic regime. Applied to the disordered transverse-field Ising model, SeTN shows that the spectral form factor is governed by the leading transfer-matrix eigenvalue, in contrast to the kicked Ising model. SeTN thus provides a novel framework for probing the disorder-driven dynamical phenomena in many-body quantum systems.
@@
Introduction.—Simulating quantum many-body systems with disorder is crucial for understanding a wide range of phenomena, including quantum chaos [gutzwillerChaosClassicalQuantum1990a, dalessioQuantumChaosEigenstate2016b, haakeQuantumSignaturesChaos2018], many-body localization (MBL) [nandkishoreManyBodyLocalizationThermalization2015c, abaninColloquiumManybodyLocalization2019b], quantum spin glass [brayReplicaTheoryQuantum1980, sachdevGaplessSpinfluidGround1993], and criticality [fisherCriticalBehaviorRandom1995, pichCriticalBehaviorGriffithsMcCoy1998]. Recently, the chaotic dynamics of disordered systems predicted by random matrix theory (RMT) has attracted growing interest, driven by advances in experimental techniques [liEmergentUniversalQuench2024, dongMeasuringSpectralForm2025].
Analytical approaches to the disorder-averaged dynamics typically map them onto classical statistical models [chanSolutionMinimalModel2018c, bertiniExactSpectralForm2018, vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019]. This mapping relies on Haar averaging [vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019] or self-duality [chanSolutionMinimalModel2018c, bertiniExactSpectralForm2018] techniques and is primarily applicable to random circuits and Floquet systems. However, it cannot be directly applied to generic time-independent disordered Hamiltonians that usually lack self-duality. Numerical methods for studying disordered dynamics mainly include exact diagonalization (ED) and tensor network (TN) techniques. ED is limited to modest system sizes, e.g., $L\lesssim 24$ for the disordered XXZ chain [pietracaprinaShiftinvertDiagonalizationLarge2018a, sierantPolynomiallyFilteredExact2020a]. While TN methods can handle larger systems, they are mostly limited to the MBL regime [doggenManybodyLocalizationDelocalization2018a, chandaTimeDynamicsMatrix2020], as extending them to the ergodic phase remains challenging due to volume-law entanglement growth.
@@
In this letter, we propose the statistics-encoded tensor network (SeTN), a general approach for studying disordered many-body systems with time-independent Hamiltonians. In SeTN, we encode disorder into an auxiliary layer and perform averaging separately (see Fig. 1(b)). By analyzing the decay rate of the singular values, we derive a criterion for the temporal resolution required for efficient disorder encoding: $n\gg\alpha^{2}t^{2}$, where $n$ is the number of time steps, $\alpha$ the disorder strength, and $t$ the evolution duration.
@@
Our approach offers two key advantages: encoding is more efficient for weak disorder, a regime challenging for conventional tensor networks, and it restores spatial translation invariance, enabling a well-defined transfer matrix that allows studying dynamics under the thermodynamic limit and naturally integrates with Krylov iteration[stewartKrylovSchurAlgorithmLarge2002] and density matrix renormalization group (DMRG) methods [whiteDensityMatrixFormulation1992a, whiteDensitymatrixAlgorithmsQuantum1993]. As a demonstration, we investigate the spectral form factor (SFF) of the disordered transverse-field Ising model in its chaotic regime, revealing a prolonged pre-RMT transient dominated by a single, non-degenerate leading eigenvalue of the transfer matrix. This pathway to RMT behavior contrasts with Floquet models such as the kicked Ising model, where chaotic spectral statistics and eigenvalue degeneracies appear from the outset [chanSolutionMinimalModel2018c].
@@
Statistics-encoded tensor network approach.—We consider a quantum many-body system with local disorder, described by the Hamiltonian $\mathcal{H}[\boldsymbol{h}]=H+\sum_{i=1}^{L}h_{i}H_{i}$, where $H$ is the disorder-free part, $\{H_{i}\}$ are local operators acting on site $i$, $\boldsymbol{h}=\{h_{i}\}$ is a disorder realization, and $L$ is the system size. Throughout this work we set $\hbar=1$. To simulate the real-time dynamics, we apply a first-order Trotter decomposition with a small time step $\tau\ll 1$, which gives the evolution operator $U(\tau)=\mathrm{e}^{-\mathrm{i}\tau H}\mathrm{e}^{-\mathrm{i}\tau\sum_{i}h_{i}H_{i}}+O(\tau^{2})$. This decomposition enables a TN representation of the full evolution operator (see Fig. 1(a)). For simplicity, we focus on the widely studied case where $H=\sum_{i}H_{i,i+1}$ includes only nearest-neighbor interactions and $H_{i}=\sigma^{z}_{i}$ is the Pauli-$z$ operator on site $i$. Generalizations to longer-range interactions or different disorder couplings are straightforward [supplementaryMaterial].
@@
where $j_{n}^{[i]}$ refers to site $i$ and time step $n$, with repeated indices summed over implicitly. This can be achieved via singular value decomposition (SVD) or spectral decomposition. At the same time, we factor out the diagonal disorder term using a delta tensor: $\langle j_{n}^{[i]}|\mathrm{e}^{-\mathrm{i}\tau\sigma^{z}_{i}}|j_{n+1}^{[i]}\rangle=v_{l_{n}^{[i]}}\delta_{l_{n}^{[i]}j_{n}^{[i]}j_{n+1}^{[i]}}$, where $\delta_{ijk}$ is the three-order Kronecker tensor (equal to $1$ when $i=j=k$ and 0 otherwise), and $\boldsymbol{v}=[\mathrm{e}^{-\mathrm{i}\tau h_{i}},\mathrm{e}^{\mathrm{i}\tau h_{i}}]$ encodes the effect of local disorder. Throughout this work, we use the indices $j,k$ and $l$ to denote the time, space, and layer directions, respectively, as indicated by the coordinate system in Fig. 1. By combining these decompositions, we define a local tensor $W$ as
@@
We study disorder-averaged dynamics, where many observables such as expectation values, the Rényi entropy, and the SFF admit a two-layer tensor-network representation, obtained by contracting the forward and backward evolutions $U$ and $U^{\dagger}$ (dark/light blue in Fig. 1(b)). The effect of disorder is encoded in a statistics layer (orange region), which after averaging renders all spatial sites equivalent. This naturally motivates the definition of a transfer matrix $\mathrm{T}$ that captures the combined action of unitary dynamics and disorder averaging across a single spatial slice,
@@
where $P(h)$ denotes the probability density function of the disorder variable $h$. We give explicit formulas for both uniform $\mathcal{U}(-\alpha,\alpha)$ and Gaussian $\mathcal{N}(0,\sigma^{2})$ disorder. While a similar expression for the Gaussian case was derived for KIM [bertiniExactSpectralForm2018], our result provides a more general formulation that applies to arbitrary systems with local disorder. This transfer matrix formulation provides a systematic and efficient framework for evaluating disorder-averaged quantities, particularly in the thermodynamic limit, and plays a central role in SeTN approach.
While the transfer matrix is a powerful tool, its implementation is limited by the exponential growth of the matrix $\mathrm{O}$, whose dimension $2^{n}\times 2^{n}$. We overcome this by a two-step SeTN compression shown in Fig. 1(c). First, each Se-Layer realization is written as a matrix product operator (MPO) with bond dimension 1; summing over $M$ such realizations via MPO addition [crosswhiteFiniteAutomataCaching2008] yields an MPO of bond dimension $M$ encoding the disorder average. Second, we iteratively compress this MPO using SVD. At $n=1$ (pink box in Fig. 1(c)), with singular values $S^{(1)}$ below a fixed threshold discarded. The truncated tensors are contracted with $A^{(2)}$ and decomposed again (blue box in Fig. 1(c)), and the process continues site by site to the final layer. Notely, since each step involves only local operations and the tensors $A^{(n)}$ are identically constructed via MPO addition from individual realizations, the overall memory requirement reduces to $O(M)$ [supplementaryMaterial], compared to the naive $O(nM^{2})$ scaling. This significantly enhances computational efficiency and enables averaging over a large number of realizations ($M\gg 1$) in practice.
Figure 2: (Color online) (a) Blue dots: squared relative singular values from SeTN compression under various parameters (Num.). Dashed lines: analytic predictions from perturbation theory (Ana.), with coefficients fitted to numerical data. (b) Blue dots: first seven coefficients $c_{i}$ obtained by fitting the data in (a). These are further fitted to exponential (green dashed), Gaussian (purple dashed), and gamma-like (red dot-dashed) functions. (c) Spectral form factor of the disordered TFIM with system size $L=4$, computed using four methods: exact diagonalizing 1000 disorder realizations (avg #1000, green dotted), numerical integration (Exact, blue dashed), SeTN (red solid), and perturbative results (Pert., gray dot-dashed). The mean deviation between SeTN and the exact result is approximately 0.021. Parameters: $J=b=1$, Trotter step $\tau=0.005$, averaging over $M=10^{6}$ realizations; singular values below $10^{-10}$ are truncated.
Properties of Se-Layer and SeTN effectiveness.—The previous section outlined the implementation of SeTN. A natural question is whether disorder averaging can be faithfully encoded into the TN representation. This effectiveness crucially depends on the behavior of the singular value spectrum $S^{(n)}$, illustrated in Fig. 1(b): only when the singular values decay rapidly, ideally exponentially, can the SeTN accurately compress the averaged transfer matrix with a low bond dimension. To analyze this, we derive an analytic expression for the spectrum by solving the recursive structure in SeTN compression perturbatively [supplementaryMaterial]:
@@
Scaling behavior of spectral form factor.—The preceding analysis shows that SeTN is more efficient for weak disorder, as singular values decay faster. In this regime, the system is typically driven into chaos, which is challenging for conventional tensor networks. To illustrate SeTN’s ability to handle chaotic dynamics, we investigate the spectral form factor (SFF) of the disordered transverse-field Ising model (disordered TFIM) in its chaotic regime.
@@
Here, we study the SFF in the disordered TFIM whose Hamiltonian reads, $\mathcal{H}[\boldsymbol{h}]=J\sum_{i=1}^{L-1}\sigma^{z}_{i}\sigma^{z}_{i+1}+b\sum_{i}^{L}\sigma^{x}_{i}+\sum_{i}h_{i}\sigma^{z}_{i}$, where $J$ is the coupling strength, $b$ is the magnetic field, and $h_{i}$ are independent random variables drawn from a uniform distribution in the range $\alpha,\alpha]$. Setting $J=b=1$ and $\alpha=0.5$ where the system is chaotic [supplementaryMaterial], we compute the SFF via ED for $L=9$ to $16$ averaging $2440$ independent disorder realizations for each size. The results clearly display the onset of the linear ramp as shown in Fig. 3(a).
@@
A key advantage of SeTN is its ability to restore spatial translational invariance broken by disorder, thereby enabling a well-defined transfer matrix and access to the thermodynamic limit. In Fig. 3(b), we show the largest eigenvalues of the transfer matrix, computed by Krylov iteration (blue crosses) and non-Hermitian DMRG (red circles) [supplementaryMaterial]. For reference, we fit the finite-size ED data in Fig. 3(a) to the form $K(t)=\lambda(t)^{L}$, with the extracted $\lambda(t)$ plotted as black lines. The excellent agreement between the largest eigenvalues and $\lambda(t)$ demonstrates that, within the accessible time window, the SFF is indeed governed by the non-degenerate leading eigenvalues of the transfer matrix. This behavior contrasts with Floquet models such as the kicked Ising model [chanSolutionMinimalModel2018c], where the SFF is controlled by a highly degenerate set of leading eigenvalues from the outset.
@@
Within the accessible time window the dynamics are governed by the non-degenerate leading eigenvalue, which is size-independent, cannot explain the size-dependent onset of the RMT ramp (the Thouless time, black dashed in Fig. 3(a)). This raises the question of how the system crosses over from the leading-eigenvalue regime to the ramp regime. Direct evaluation of the eigenvalues at the Thouless time is hindered by rapid bond growth in Krylov iteration and the lack of convergence in non-Hermitian DMRG. Inspired by the observed near degeneracy of the leading eigenvalues where the confidence interval (shaded region) expands, we conjecture that around the Thouless time, the leading eigenvalues become nearly degenerate, with the degree of near-degeneracy increasing in time. Such increasing near degeneracy implies that at finite sizes, multiple eigenvalues contribute comparably, producing the ramp, whereas in the thermodynamic limit the single largest eigenvalue eventually dominates. This scenario naturally explains both the emergence of the RMT ramp and the size-dependence of the Thouless time, as further illustrated by a toy-model analysis [supplementaryMaterial]. These insights highlight the potential of SeTN-defined transfer matrices in capturing the crossover to RMT spectral statistics in disordered time-independent quantum systems.
Figure 3: (Color online) (a) SFF of the disordered TFIM with sizes $L=9$ to $16$ (top to bottom), computed by ED using 2440 independent disorder realizations for each size. (b) Magnitudes of the leading and subleading eigenvalues of the SeTN-derived transfer matrix as functions of time, obtained using the Krylov method (blue crosses) and non-Hermitian DMRG (red circles). Black lines represent the fitted coefficient extracted by fitting the data in panel (a) to the form $\lambda(t)^{L}$, and the shaded region indicates the 80% confidence interval. Parameters: Trotter step $\tau=0.05$, $J=b=1,\alpha=0.5$.

### v2

@@ Abstract
Simulating the dynamics of quantum many-body systems with disorder is a fundamental challenge. In this work, we propose a general approach—the statistics-encoded tensor network (SeTN)—to study such systems. By encoding disorder into an auxiliary layer and averaging separately, SeTN restores translational invariance, enabling a well-defined transfer matrix formulation. We derive a universal criterion, $n\gg\alpha^{2}t^{2}$, linking discretization $n$, disorder strength $\alpha$, and evolution duration $t$. This sets the resolution required for faithful disorder averaging and shows that encoding is most efficient in the weak-disorder, typically chaotic regime. Applied to the disordered transverse-field Ising model, SeTN shows that, over the numerically accessible time window, the spectral form factor is governed by the leading transfer-matrix eigenvalue, in contrast to the kicked Ising model. SeTN thus provides a novel framework for probing the disorder-driven dynamical phenomena in many-body quantum systems.
I Introduction
@@
Simulating quantum many-body systems with disorder is crucial for understanding a wide range of phenomena, including quantum chaos [gutzwillerChaosClassicalQuantum1990a, dalessioQuantumChaosEigenstate2016b, haakeQuantumSignaturesChaos2018], many-body localization (MBL) [nandkishoreManyBodyLocalizationThermalization2015c, abaninColloquiumManybodyLocalization2019b, sierantManybodyLocalizationAge2025], quantum spin glass [brayReplicaTheoryQuantum1980, sachdevGaplessSpinfluidGround1993], and criticality [fisherCriticalBehaviorRandom1995, pichCriticalBehaviorGriffithsMcCoy1998]. Recently, the chaotic dynamics of disordered systems predicted by random matrix theory (RMT) has attracted growing interest, driven by advances in experimental techniques [liEmergentUniversalQuench2024, dongMeasuringSpectralForm2025].
Analytical progress on disorder-averaged dynamics has primarily been achieved in random quantum circuits and dual-unitary models, where Haar averaging or self-duality enables exact transfer-matrix treatments [chanSolutionMinimalModel2018c, bertiniExactSpectralForm2018, vonkeyserlingkOperatorHydrodynamicsOTOCs2018a, zhouEmergentStatisticalMechanics2019, nahumOperatorSpreadingRandom2018c, sonnerCharacterizingManybodyLocalization2022, riddellStructuralStabilityHypothesis2024, bertiniExactlySolvableManybody2025a]. However, these approaches cannot be directly extended to generic time-independent Hamiltonians, where self-duality is absent and the transfer-matrix dimension grows exponentially, hindering practical implementation. Numerically, exact diagonalization (ED) is restricted to small system sizes [pietracaprinaShiftinvertDiagonalizationLarge2018a, sierantPolynomiallyFilteredExact2020a], while conventional tensor-network methods and strong-disorder renormalization approaches [doggenManybodyLocalizationDelocalization2018a, chandaTimeDynamicsMatrix2020, hikiharaNumericalRenormalizationgroupStudy1999, goldsboroughSelfassemblingTensorNetworks2014a] treat each disorder realization independently and are applicable mostly to the deep MBL regime. Another widely used route in this regime, quantum-parallel method [alvarezQuantumParallelismTool2008a, paredesExploitingQuantumParallelism2005c, andraschkoPurificationManyBodyLocalization2014, hubigTimedependentStudyDisordered2019a], embeds disorder realizations taking only discrete values into an auxiliary system and evolves them in parallel. However, when the disorder follows a continuous distribution, as is typical in studies of chaotic dynamics, the method requires an auxiliary space of infinite dimension.
@@
In this letter, we propose the statistics-encoded tensor network (SeTN), a general approach for studying disordered many-body systems with time-independent Hamiltonians. In SeTN, we encode disorder into an auxiliary layer on which the average is taken independently at each site (Fig. 1(b)). By analyzing the decay rate of the singular values, we derive a universal criterion for the temporal resolution required for efficient disorder encoding: $n\gg\alpha^{2}t^{2}$, where $n$ is the number of time steps, $\alpha$ the disorder strength, and $t$ the evolution duration. This criterion shows that continuous disorder distributions can be efficiently encoded, particularly in the weak-disorder regime, making disorder-averaged transfer matrices numerically feasible for generic Hamiltonians.
@@
After averaging, SeTN restores spatial translation invariance, yielding a compact MPO transfer matrix that can be directly analyzed using Krylov iteration [stewartKrylovSchurAlgorithmLarge2002] and density matrix renormalization group (DMRG) methods [whiteDensityMatrixFormulation1992a, whiteDensitymatrixAlgorithmsQuantum1993]. As a demonstration, we investigate the spectral form factor (SFF) of the disordered transverse-field Ising model in its chaotic regime, revealing, within the numerically accessible regime, a prolonged pre-RMT transient dominated by a single, non-degenerate leading eigenvalue of the transfer matrix. This pathway to RMT behavior contrasts with Floquet models such as the kicked Ising model, where chaotic spectral statistics and eigenvalue degeneracies appear from the outset [chanSolutionMinimalModel2018c].
II Statistics-encoded tensor network approach
@@
We consider a quantum many-body system with local disorder, described by the Hamiltonian $\mathcal{H}[\boldsymbol{h}]=H+\sum_{i=1}^{L}h_{i}H_{i}$, where $H$ is the disorder-free part, $\{H_{i}\}$ are local operators acting on site $i$, $\{h_{i}\}$ are disorder variables which are independent and identically distributed, and $L$ is the system size. Throughout this work we set $\hbar=1$. To simulate the real-time dynamics, we apply a first-order Trotter decomposition with a small time step $\tau\ll 1$, which gives the evolution operator $U(\tau)=\mathrm{e}^{-\mathrm{i}\tau H}\mathrm{e}^{-\mathrm{i}\tau\sum_{i}h_{i}H_{i}}+O(\tau^{2})$. This decomposition enables a TN representation of the full evolution operator (see Fig. 1(a)). For simplicity, we focus on the widely studied case where $H=\sum_{i}H_{i,i+1}$ includes only nearest-neighbor interactions and $H_{i}=\sigma^{z}_{i}$ is the Pauli-$z$ operator on site $i$. Generalizations to longer-range interactions or different disorder couplings are straightforward [supplementaryMaterial].
@@
where $j_{n}^{[i]}$ refers to site $i$ and time step $n$, with repeated indices summed over implicitly. This can be achieved via singular value decomposition (SVD) or spectral decomposition. At the same time, we factor out the diagonal disorder term using a delta tensor: $\langle j_{n}^{[i]}|\mathrm{e}^{-\mathrm{i}\tau\sigma^{z}_{i}}|j_{n+1}^{[i]}\rangle=v_{l_{n}^{[i]}}\delta_{l_{n}^{[i]}j_{n}^{[i]}j_{n+1}^{[i]}}$, where $\delta_{ijk}$ is the three-order Kronecker tensor (equal to $1$ when $i=j=k$ and 0 otherwise), and $v_{l}=\mathrm{e}^{-\mathrm{i}\tau h_{i}l}$ with $l\in\{1,-1\}$ encodes the effect of local disorder. Throughout this work, we use the indices $j,k$ and $l$ to denote the time, space, and layer directions, respectively, as indicated by the coordinate system in Fig. 1. By combining these decompositions, we define a local tensor $W$ as
@@
Figure 2: (Color online) (a) Blue dots: squared relative singular values from SeTN compression under various parameters (Num.). Dashed lines: analytic predictions from perturbation theory (Ana.), with coefficients fitted to numerical data. (b) Blue dots: first seven coefficients $c_{i}$ obtained by fitting the data in (a). These are further fitted to exponential (green dashed), Gaussian (purple dashed), and gamma-like (red dot-dashed) functions. (c) Spectral form factor of the disordered TFIM with system size $L=4$, computed using four methods: exact diagonalizing 1000 disorder realizations (avg #1000, green dotted; shaded band indicates the standard error), numerical integration (Exact, blue dashed), SeTN (red solid), and perturbative results (Pert., gray dot-dashed). The mean deviation between SeTN and the exact result is approximately 0.021. Parameters: $J=b=1$, Trotter step $\tau=0.005$, averaging over $M=10^{6}$ realizations; singular values below $10^{-10}$ are truncated.
We study disorder-averaged dynamics, for which many quantities of interest including averaged operator expectation values $\langle\!\langle O(t)\rangle\!\rangle=\langle\,\operatorname{Tr}[O\,U(t)\rho(0)U^{\dagger}(t)]\,\rangle$, as well as the SFF $K(t)=\langle\,\operatorname{Tr}[U(t)\otimes U^{\dagger}(t)]\,\rangle$, naturally admit a two-layer tensor-network representation obtained by contracting the forward and backward evolutions $U$ and $U^{\dagger}$ (dark/light blue in Fig. 1(b)). The effect of disorder is encoded in a statistics layer (orange region) which, after averaging, renders all spatial sites equivalent. This structure motivates the definition of a transfer matrix $\mathrm{T}$ that captures the combined action of unitary dynamics and disorder averaging across a single spatial slice,
@@
where $P(h)$ denotes the normalized probability density function of the disorder variable $h$, i.e., $\int dh\,P(h)=1$. We give explicit formulas for both uniform $\mathcal{U}(-\alpha,\alpha)$ and Gaussian $\mathcal{N}(0,\sigma^{2})$ disorder. While a similar expression for the Gaussian case was derived in the KIM context [bertiniExactSpectralForm2018], the present formulation applies to general static Hamiltonians with arbitrary on-site disorder. Moreover, when $P(h)$ is discrete, $\mathrm{O}$ reduces exactly to the standard quantum-parallel encoding [supplementaryMaterial].
Although Eq. 4 represents the ideal analytic average, directly constructing the $2^{n}\times 2^{n}$ tensor $\mathrm{O}$ is numerically infeasible. In practice, we therefore adopt a finite-sample approximation with $M$ disorder realizations sampled from the distribution $P(h)$, and compress the resulting MPO through the two-step procedure illustrated in Fig. 1(c). First, each Se-Layer realization is written as a matrix product operator (MPO) with bond dimension 1; summing over $M$ such realizations via MPO addition [crosswhiteFiniteAutomataCaching2008] yields an MPO of bond dimension $M$ encoding the disorder average. Second, we iteratively compress this MPO using SVD. At $n=1$ (pink box in Fig. 1(c)), the singular values $S^{(1)}$ below a fixed threshold are discarded. The truncated tensors are contracted with $A^{(2)}$ and decomposed again (blue box in Fig. 1(c)), and the process continues site by site to the final layer.
Notely, since each step involves only local operations and the tensors $A^{(n)}$ are identically constructed via MPO addition from individual realizations, the overall memory requirement reduces to $O(M)$ [supplementaryMaterial], compared to the naive $O(nM^{2})$ scaling. This significantly enhances computational efficiency and enables averaging over a large number of realizations ($M\gg 1$) in practice.
III Properties of Se-Layer and SeTN effectiveness
The previous section outlined the implementation of SeTN. A natural question is whether disorder averaging can be faithfully encoded into the TN representation. This effectiveness crucially depends on the behavior of the singular value spectrum $S^{(n)}$, illustrated in Fig. 1(b): only when the singular values decay rapidly, ideally exponentially, can the SeTN accurately compress the averaged transfer matrix with a low bond dimension. To analyze this, we derive an analytic expression for the spectrum by solving the recursive structure in SeTN compression perturbatively [supplementaryMaterial]:
@@ III Properties of Se-Layer and SeTN effectiveness
Figure 3: (Color online) (a) SFF of the disordered TFIM with sizes $L=9$ to 1616 (top to bottom), computed by ED using 2440 independent disorder realizations per size. Standard errors are below 4% at all times, and are therefore not visible on the logarithmic scale. Shaded regions (i) and (ii) indicate the first and second rebounce windows. (b) Magnitudes of the leading and subleading eigenvalues of the SeTN-derived transfer matrix as functions of time, obtained using Krylov (blue crosses) and non-Hermitian DMRG (red circles). Horizontal bars (i) and (ii) match the time windows in (a), highlighting the near-degeneracy of dominant eigenvalues. Black lines represent the fitted coefficient extracted by fitting the data in panel (a) to the form $\lambda(t)^{L}$, and the shaded region indicates the 80% confidence interval. Parameters: Trotter step $\tau=0.05$, $J=b=1,\alpha=0.5$.
IV Scaling behavior of spectral form factor
The preceding analysis shows that SeTN is more efficient for weak disorder, as singular values decay faster. In this regime, the system is typically driven into chaos, which is challenging for conventional tensor networks. To illustrate SeTN’s ability to handle chaotic dynamics, we investigate the spectral form factor (SFF) of the disordered transverse-field Ising model (disordered TFIM) in its chaotic regime.
@@ IV Scaling behavior of spectral form factor
Here, we study the SFF in the disordered TFIM whose Hamiltonian reads, $\mathcal{H}[\boldsymbol{h}]=J\sum_{i=1}^{L-1}\sigma^{z}_{i}\sigma^{z}_{i+1}+b\sum_{i}^{L}\sigma^{x}_{i}+\sum_{i}h_{i}\sigma^{z}_{i}$, where $J$ is the coupling strength, $b$ is the magnetic field, and $h_{i}$ are independent random variables drawn from a uniform distribution in the range $\alpha,\alpha]$. We use open boundary conditions throughout; the results are insensitive to the choice of boundary condition. Setting $J=b=1$ and $\alpha=0.5$ where the system is chaotic [supplementaryMaterial], we compute the SFF via ED for $L=9$ to $16$ averaging $2440$ independent disorder realizations for each size. The results clearly display the onset of the linear ramp as shown in Fig. 3(a).
@@
A key advantage of SeTN is its ability to restore spatial translational invariance broken by disorder, thereby enabling a well-defined transfer matrix and access to the thermodynamic limit. In Fig. 3(b), we show the largest eigenvalues of the transfer matrix, computed by Krylov iteration (blue crosses) and non-Hermitian DMRG (red circles) [supplementaryMaterial]. For reference, we fit the finite-size ED data in Fig. 3(a) to the form $K(t)=\lambda(t)^{L}$, with the extracted $\lambda(t)$ plotted as black lines. The excellent agreement between the largest eigenvalues and $\lambda(t)$ demonstrates that, within the accessible pre-ramp time window, the SFF is indeed governed by the non-degenerate leading eigenvalues of the transfer matrix. This behavior contrasts with Floquet models such as the kicked Ising model [chanSolutionMinimalModel2018c], where the SFF is controlled by a highly degenerate set of leading eigenvalues from the outset.
Within the accessible pre-ramp time window, the dynamics are governed by the non-degenerate leading eigenvalue, which is size independent, cannot explain the size-dependent onset of the RMT ramp (the Thouless time, black dashed in Fig. 3(a)). Understanding how the system crosses over from the leading-eigenvalue regime to the ramp regime thus requires access to longer times.
@@
At present, a direct evaluation of the eigenvalues in the ramp region is hindered by rapid bond growth in Krylov iteration and the lack of convergence in non-Hermitian DMRG. Looking beyond the numerically accessible regime, the observed broadening of the confidence intervals and the near-degeneracy of the leading eigenvalues suggest a possible crossover scenario. In particular, we conjecture that around the Thouless time the leading eigenvalues become increasingly near-degenerate, such that at finite sizes multiple eigenvalues contribute comparably, producing the ramp, whereas in the thermodynamic limit the single largest eigenvalue eventually dominates. This conjectured scenario is consistent with the size dependence of the Thouless time and is further illustrated by a toy-model analysis [supplementaryMaterial].
V Summary and outlook